# Nonuniqueness of capped black holes: large and small bubbles

### v1

@@ IV.3 Nonuniqueness of spherical black holes: large and small bubble branches
Figure 7: Entropy of large and small bubble branches for $\sigma=0.8$ and $j_{\phi}=0.235$ displayed by the blue and red curves, respectively. The entropy of the Cvetič-Youm black hole with the same charges is also shown by the gray curve.
IV.4 Relation with three parameter solution Suzuki:2023nqf ; Suzuki:2024phv
Finally, we discuss the connection between the new four-parameter solution and the previously three-parameter capped black hole solution, as detailed in Refs.Suzuki:2023nqf ; Suzuki:2024phv . This three-parameter solution is derived from the four-parameter model by simply setting $\beta=0$. As shown in Fig. 8, the $\beta=0$ phase is situated near the $\tilde{\gamma}=0$ curve. It begins at the singular extremal Cvetič-Youm phase (marked as $B$ in Fig. 8) with $\tilde{\nu}=\sqrt{2}$ and ends at the zero entropy extremal capped black hole phase (marked as $C^{\prime}$ in Fig. 8) with $\tilde{\nu}=1$. Notably, within the three-parameter family, the $\beta=0$ phase does not exhibit overlaps in the $(j_{\psi},j_{\phi})$-plane, thus not displaying the non-uniqueness associated with the four-parameter family.
Figure 8: The phase of the three parameter capped black hole solution (black curve) embedded in the four parameter phase space for $\sigma=0.8$ presented in Fig. 4.
@@ V Summary
The phase diagram we have obtained may not complete since the parameter region is partially bounded by nonextremal, nonsingular phases, indicating the possible extension of the parameter region as in Ref. Suzuki:2024vzq . In the extended region, some parameters would take complex values, even if the metric and gauge potential remain real. It may also be an interesting issue to construct non-BPS black hole solutions with the DOC of other non-trivial topology, $[{\mathbb{R}}^{4}\#n{\mathbb{C}}{\mathbb{P}}^{2}\#m(S^{2}\times S^{2})]\setminus{\mathbb{B}}^{4}$ with $(n,m)\not=(1,0)$ on a timeslice.

### v2

@@ IV.3 Nonuniqueness of spherical black holes: large and small bubble branches
Figure 8: Entropy of large and small bubble branches for $n=1$, $\sigma=0.8$ and $j_{\phi}=0.235$ displayed by the blue and red curves, respectively. The entropy of the Cvetič-Youm black hole with the same charges is also shown by the gray curve.
IV.4 Relation to the three-parameter solution in Ref. Suzuki:2023nqf ; Suzuki:2024phv
Finally, we discuss the connection between the new four-parameter solution and the previous three-parameter solution, as detailed in Refs.Suzuki:2023nqf ; Suzuki:2024phv . This three-parameter solution is derived from the four-parameter solution by simply setting $\beta=0$. As shown in Fig. 9, the $\beta=0$ phase is represented by a curve starting from $A$ with $\tilde{\nu}=\sqrt{2}$ and ending at $C=C^{\prime}$ with $\tilde{\nu}=1$ in the $(j_{\psi},j_{\phi})$ and $(\phi_{\mathcal{D}},j_{\phi})$ planes. Whereas, in the $(q_{\mathcal{D}},j_{\phi})$ plane, the end point with $\tilde{\nu}=1$ appears on the $j_{\phi}=0.240\ldots$ line between $C$ and $C^{\prime}$. Because the $\beta=0$ phase does not exhibit multiple branches in the $(j_{\psi},j_{\phi})$-plane, the non-uniqueness of the capped black hole, discussed in the previous subsection, cannot be confirmed within three parameter-parameter solutions.
Figure 9: The phase of the three parameter capped black hole solution (black curve) embedded in the four parameter phase space for $n=1$ and $\sigma=0.8$ presented in Fig. 4.
@@ V Summary
We have found that, as long as keeping all parameters real, the $Q/M={\rm constant}$ phase in the $n=1$ case is not complete, but it also requires the solution obtained from the $Q/M={\rm constant}$ phase in the $n=-1$ case via the transformation $(t,\psi)\to(-t,-\psi)$. The complete phase space may be expressed by using only the $n=1$ solution if one can extend the parameter region to include certain complex parameters, as preformed in Ref. Suzuki:2024vzq . This extension deserves further investigation in our future work. It may also be an interesting issue to construct non-BPS black hole solutions with the DOC of other non-trivial topology, $[{\mathbb{R}}^{4}\#n{\mathbb{C}}{\mathbb{P}}^{2}\#m(S^{2}\times S^{2})]\setminus{\mathbb{B}}^{4}$ with $(n,m)\not=(1,0)$ on a timeslice. This extension will be the focus of our subsequent studies.